Black Scholes Option Pricing Model

Stock Price = 75 Strike price = 70 Risk Free rate - 4% Standard deviation = 15% 5 months remaining

Calculate call & Put and show work please
Answer #1

d1=[ln(S/X) + (R – d + σ^2 / 2) × t] / (σ ×t^0.5) =[Ln(75/70)+(4%-0%+15%^2/2)*5/12]/(15%*5/12^0.5) =0.9331

N(d1) using NORMSDIST function of excel =NORMSDIST(0.9331) =0.8246

d2 =d1-σ*t^0.5 =0.9331-15%*5/12^0.5=0.8363
N(d2) using NORMSDIST function of excel =NORMSDIST(0.8363) =0.7985

Call Option =C = S * N(d1) – X × e–Rt × N(d2) =75*0.9331-70*EXP(-4%*5/12)*0.7985 =6.69

P=C0-S0+X*EXP-rt =6.69-75+70*EXP(-4%*5/12) =0.53